OpenStudy (anonymous):

Two consecutive odd integers have a sum of 48. What is the largest of the two integers? A. 23 B. 24 C. 25 D. 26

OpenStudy (anonymous):

1st number = x 2nd number = x + 2 x + (x + 2) = 48 2x + 2 = 48 2x = 46 x = 23 1st number = x = 23 2nd number = x + 2 = 25 the largest of the two integers is 25. C. 25